(a) 2(2x + 1)
(b) 4(3x-5)
(c) (7a +2)3
(d) 3m(2-n)
(e) 4m (3+5n)

Answer by checkley71(8403)

a) 2(2x+1)=4x+2
b) 4(3x+5)=12x+20
c) (7a+2)3=21a+6
d) 3m(2-n)=6m-3mn
e) 4m(3+5n)=12m+20mn